Solve the following equation of first order but not of first degree y=2xp-p²


1
Expert's answer
2021-05-21T07:08:45-0400

"\text{Given, the equation} y=2xp-p².\newline
\text{Differentiate wrt} x.\newline
p=2p-2pr\newline
\text{where}p=\frac{dy}{dx} \text{and }\frac{dp}{dx}.\newline
1=2-2r\newline
r=\frac{1}{2}\newline
\frac{dp}{dx}=\frac{1}{2}\newline
dp=\frac{1}{2}dx\newline
\text{Integrate wrt}x,\newline
p=\frac{x}{2}+C\newline
\frac{dy}{dx}=\frac{x}{2}+C\newline
dy=(\frac{x}{2}+C)dx\newline
\text{Integrate wrt} x,\newline
y=\frac{x^2}{4}+Cx+D"